OpenStudy (anonymous):

Write the equation of a line in point-slope form, when the coordinate is (-1,2) ?

OpenStudy (anonymous):

y-2=m (x-(-1)) y=m(x+1) +2 y=mx + (2+m)

OpenStudy (anonymous):

what is that in general form?

OpenStudy (anonymous):

general form: mx-y -(2+m)=0